OpenStudy (anonymous):

Identify the graph: f(x) = (x - 2)2 - 3

OpenStudy (anonymous):

@ cosmichaotic

OpenStudy (tkhunny):

This is the Vertex Form of a parabola. Vertex: (2,-3) y-Intercept (0,1) Lots of other things, but those are the most obvious.

OpenStudy (anonymous):

@tkhunny : false. f(x) = (x-2)2-3 f(x) = 2x-4-3 f(x) = 2x-7

OpenStudy (anonymous):

f(x) = -(x - 9)^2 + 8? @Cosmichaotic or something

OpenStudy (tkhunny):

Generally, if one means 2 * some quantity, the "2" appears first. Often, if one means some quantity raised to the second power, the 2 appears last. It is a matter of opinion until the OP clarifies it.

OpenStudy (tkhunny):

@alyygirl This one is EXACTLY like the last one. You tell us about it, please.

OpenStudy (unklerhaukus):